Can you please explain to me how you get the answer to this as well as how to do this .
If f(x)=2x-6 and g(x)=x-2x*2 , Find each value
f(2)
g(-1/3)

Answer by ankor@dixie-net.com(22740)

If f(x)=2x-6 and g(x)=x-2x*2 , Find each value
f(2)
x = 2 (replace x with 2)
f(2) = 2(2) - 6
f(2) = 4 - 6
f(2) = -2
:
g(-1/3)
x =-1/3 (replace x with -1/3)
g(-1/3) = (-1/3) - 2(-1/3)^2
g(-1/3) = (-1/3) - 2(1/9)
g(-1/3) = (-1/3) - (2/9)
g(-1/3) = (-3/9) - (2/9)
g(-1/3) = -5/9